OpenStudy (anonymous):

If x varies inversely as y, and x = 11 when y = 15, find x when y = 3. X=

OpenStudy (anonymous):

First it is key to understand what the term "inversely" means exactly. x varies inversely to y implies y=1/x or x=1/y (they are equivalent). You may multiply the equation by a constant and this does not affect that condition.Does this help you to answer your question?

OpenStudy (hba):



OpenStudy (anonymous):

Thank you both for your help.